OpenStudy (anonymous):

An equation of a circle is x^2 + y^2 + 10x + 6y + 18 = 0. Part 1: Show all your work in determining the center and radius of this circle. (3 points) Part 2: In complete sentences, explain the procedure used. (3 points)

OpenStudy (anonymous):

I have no clue how to do this one :(

OpenStudy (anonymous):

\[x^2 + y2 + 10x + 6y + 18 = 0\] complete the square twice first rewrite as \[x^2+10x+y^2+6y=-18\] do you know how to complete a square?

OpenStudy (anonymous):

what is half of 10?

OpenStudy (anonymous):

I don't remember... Can you remind me?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

good, what is \(5^2\) ?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

center is (-5,-3) and radius is sqrt(15)

OpenStudy (anonymous):

ok so the first thing we are going to write instead of \[x^2+10x+y^2+6y=-18\] is \[(x+5)^2+y^2+6y=-18+25\] or \[(x+5)^2+y^2+6y=7\]

OpenStudy (anonymous):

now what is half of 6?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

btw answer given above is wrong

OpenStudy (anonymous):

good, okay, and again, what is \(3^2\) ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

@amistre64 can you help me finish this? @satellite73 left...

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

ok so the first thing we are going to write instead of \(x^2+10x+y2+6y=-18\) is \((x+5)^2+y2+6y=-18+25\) or \((x+5)^2+y2+6y+9=7+9\)

ganeshie8 (ganeshie8):

ok so the first thing we are going to write instead of \(x^2+10x+y^2+6y=-18\) is \((x+5)^2+y^2+6y=-18+25\) or \((x+5)^2+y^2+6y+9=7+9\) or \((x+5)^2+(y+3)^2=16\)

ganeshie8 (ganeshie8):

now, u knw the standard form of equation of circle wid center (h, k) and radius r ?

OpenStudy (anonymous):

@ganeshie8 I got the center part, but what's the radius?

ganeshie8 (ganeshie8):

oh nice :) wat do u get for center ?

OpenStudy (anonymous):

(-5,-3)

ganeshie8 (ganeshie8):

thats correct !

OpenStudy (anonymous):

but how do I get the radius?

ganeshie8 (ganeshie8):

\((x-h)^2+(y-k)^2=r^2\) center = \((h, k)\) radius = \(r\)

OpenStudy (anonymous):

so r=4?

ganeshie8 (ganeshie8):

just observe that 16 can be written as 4^2

ganeshie8 (ganeshie8):

Yes ! r = 4

OpenStudy (anonymous):

Thanks soooooooooooooooo much!!!!

ganeshie8 (ganeshie8):

np :D

ganeshie8 (ganeshie8):

and plz medal satellite, not me :)